OpenStudy (anonymous):

solve 2sqrtx7 – 10 = –8

OpenStudy (anonymous):

\[2\sqrt{x +7 } -10 =-8\]

OpenStudy (anonymous):

You wanna get the square root alone on one side first

OpenStudy (anonymous):

So i square both sides?

OpenStudy (anonymous):

No, you need to get it alone first, so add 10 to both sides and then divide both sides by 2

OpenStudy (anonymous):

so sqrt x-+7= 1?

OpenStudy (anonymous):

sqrt(x+7)=1 yes now square both sides